EVA Report – April 29th

Sol 7 Crew 193 EVA 13 & 14 Summary
Earth-Date 29APR2018

EVA 13

EVA Gold Crew Members – Gold000 (CO Patil-Sabale) / Gold 003 (CE Attig) / Gold005 (CGA Shear)

EVA Objectives – Summary
1. Explored White Rock Canyon, and Kissing Camel Ridge. Found a few potential shelter sites, and a small pond of muddy water!
2. Repeated GPS recording experiment.

3. Continue to explore the crew biometrics study while on EVA. Raw data here
4. Continuing visor fogging study. Raw data here

Issues/Mitigation

  • Was super windy. No major issues

ATVs/Rovers Used

  • ATV 300
  • Honda ATV
  • ATV #350.3

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glove precautions – Yes

EVA 14

EVA Gold Crew Members – Gold001 (XO Campbell) / Gold 004 (HSO Pandya) /

EVA Objectives – Summary

  1. Explore Toothy Ridge Area and Copernicus hwy
  2. Continue to explore the crew biometrics study while on EVA / Raw data here
  3. Continued search for emergency shelters
  4. Continuing visor fogging study / Raw data here

Issues/Mitigation

  • Extreme winds.

ATVs/Rovers Used
Honda ATV
ATV #350.3

Safety Measures & Preparation

  • Sunscreen/Sun protection – Yes/Medical Exemptions for UV eye-protection provided
  • Adequate hydration – No
  • Adequate protection – Flightsuit/footwear/visor/glov e precautions – Yes

Warm Regards!
Commander, (Gold 00), Crew #193

Mars Desert Research Station